OpenStudy (anonymous):

1+3(5-17)divided by 2-6

OpenStudy (anonymous):

8.75

OpenStudy (anonymous):

1+3(5-17) 1+3(-12) 1+(-36) -35 2-6 = -4 so -35/-4 = 8.75

OpenStudy (anonymous):

\[\frac{1+3(5-17)}{2-6} =\frac{1+3(-12)}{-4}=\frac{1-36}{-4}=\frac{-35}{-4}=8.75\]